OpenStudy (anonymous):

what is the standard form equation of the ellipse with vertices at (0,6) and (0,-6) and co vertices at (4,0) and (-4,0)?

OpenStudy (aum):

What is the length of the semi-major axis? What is the length of the semi-minor axis?

OpenStudy (anonymous):

um thats all the information it gives....

OpenStudy (aum):

Yeah, from the given information you can figure out the lengths of semi-major and semi-minor axes. 

OpenStudy (anonymous):

okay so would the equation be x^2/16+y^2/36=1

OpenStudy (aum):

Since the major axis is vertical, along the y-axis, the equation of the ellipse is: \[\large \frac{x^2}{b^2}+\frac{y^2}{a^2} = 1 \]where a is the length of the semi-major axis and b is the length of the semi-minor axis.

OpenStudy (aum):

Correct.

OpenStudy (anonymous):

wow thank youuuu

OpenStudy (aum):

You are welcome.